Has anybody experienced problems about the profile changes in AD9914? It seems that when more than 1 bit of the profile select bits change simultaneously the output sometimes gets messy. I started to investigate the problem however I wanted to ask to the community if anyone has experienced a similar problem before hand as it seems to take time to fully understand it.
I tried changing the profiles as indicated below and found the corresponding results so far :
state change between 001 and 011 is PASS
state change between 010 and 011 is PASS
state change between 100 and 101 is PASS
state change between 001 and 010 is FAIL
state change between 100 and 011 is FAIL
I will try the other posibilities also.
Different profile change speeds from 96ns to 1us for the failing conditions are tested however results don't change.
Using a single profile and just updating its contents on the fly I managed to implement every frequency change I wanted. I just tried this to see if I could get a clue about the Sync_Clk problem but there seems to be none.
The profile selects pins are driven by a buffer and observed them by an oscilloscope and there isn't any problem about the drive lines.
I tried another board and the results are same.( Both of the boards are our custom board)
May there be any problem about 3.5GHz clocking rate of the core? Because this maximizes the frequency of sync_clk with which internal logic runs. I will try lower sysclk rates in the upcoming days.
Sync_Clk is not used.
Any help is highly appreciated.
Message was edited by: oytun kurt I tried all 12 profile change possibilities with a single bit change and everything works fine. I tried lower sysclk rate that is 2.8GHz and same problem continues with double or triple profile bit changes. Same problem has been generated with the eval board. I programmed Profile1 with 400MHz and Profile2 with 120M. I recorded the output with an oscilloscope by triggering with the profile change from 2 to 1. For 700us after the profile change the output is nearly garbage. Then it settles to 400MHz. I observed the frequency contents of the garbage section by using the oscilloscope's spectrum option and found out that there are two main frequency components; first at 1.25GHz and the second at 1.75GHz. I don't know what to do more to figure out the problem and waiting for your help. Regards, Oytun